United Claim Processors provides claims processing services for major insurance companies. The Winnipeg location processes all health benefit claims for various insurance benefit providers. Claims are received electronically or by mail in the Initial Review Department. Employees in this department sort claim into simple and complex categories and note any missing documentation that may be required. Simple claims are then passed on to the Processing Department where skilled claims processors match information from the claim with each individual’s accounts in the database and then estimate the value of the benefit payment to be paid. Complex claims go to a separate processing center for further examination. The next stage in the simple claims process is the Approval and Payment Department where final approvals and payments to claimants are initiated and completed.

During March, the Initial Review Department transferred 5,000 simple claims to the Processing Department. At the end of March, 800 claims remained in the Initial Review Department that was 80% complete with respect to conversion. The cost per equivalent unit for the month was $1.50 per claim processed and total costs incurred in the Initial Review Department during the month totalled $8,460. (Note there are only conversion costs in this process; no raw materials are added at any point.)

Required:

Prepare the cost reconciliation portion of the Initial Review Department’s production report for March.

Conversion EU Total Cost Cost accounted for as follow: Transferred to Packaging Dept: Work in process, March 31 Conversion Total cost accounted for 0
Answer #1
Total cost Conversion EU
Cost accounted for as follows:
Transferred to Packaging Dept: 7500 5000
Work in process, March 31:
Conversion 960 640
Total cost accounted for 8460
Workings:
Work in process, March 31 EU = 800*80%=640
Transferred to Packaging Dept cost=5000*1.50=$7500
Work in process, March 31 cost = 640*1.50= $960